GW014

PI: Wiik
Title: 15 AGNs

Session: Oct 1996 (session 3)
Frequency: 1.3 cm
Mode: E
Obs time (UT): 293 EU 05:00(19/10)-05:00(20/10)
Obs time (UT): 293 US 05:00(19/10)-05:00(20/10)
Stations: EbJbMhNtOnSh
Processor: Soc
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Metsahovi (Kirsi Karlamaa, Ari Mujunen, Kaj Wiik at Mon Oct 21 03:47:25 EDT 1996) Minor failures
Power outage at 293 09:52--10:15. Two suspicious head position log lines at 293 07:13:1145/pass/10,,auto,-220.0,,54.7,,274.7, and 293 09:13:1139/pass/18,,auto,-110.0,,54.7,,164.7,. In both cases the head position should have stayed the same, i.e. the FS has probably not attempted to move the head but for some unknown reason it gives a wrong LVDT readout (or the head really has moved to a wrong position on its own...) Rainy weather. Tapes to Socorro are being shipped today on Monday as agreed in TWG; Bonn will follow on Tuesday tomorrow.

Effelsberg (D.Graham at Wed Oct 23 09:08:35 EDT 1996) Minor failures
The PI wrote his own Tsys procedures, to be used instead of the standard ones. These did not work: now the PI has no Tsys information at all. Otherwise uneventful.

Onsala (J.Conway, A. Polatidis at Thu Oct 24 07:22:26 EDT 1996) Minor failures
The Tsys and TPI procedures provided were not standard for Mark III. We have edited out both procedures and Tsys and TPI measurements are reliable after 1100 UT. Missed three scans between ~1100-1200 UT when correcting the Tsys/Tpi procedure. We also lost a scan at 151500-152800 due to an operator error. No other problems known. Rain during parts of the experiment.

Noto (Corrado Trigilio at Wed Oct 30 05:26:58 EST 1996) Minor failures
Tsys procedures have been previously edited (because we have a VLBA terminal) and checked, as required from the PI. They worked fine! At the beginning of many scans, the telescope was still slewing, so Tsys is done offsource. Clear sky during the whole experiment.

Jodrell (Alastair Gunn at Thu Nov 7 06:27:26 EST 1996) Minor failures
PI supplied non-standard Tsys procedure which did not work so no Tsys information was recorded. Due to staff shortages ay Jodrell it is not possible for us to do detailed checks on all schedules and procedures, particularly when complicated procedures are required. PIs should ensure that non-standard procedures will work or supply appropriate details for changes. Time allowed for source changes was insufficient probably affecting ALL source changes. Major problems with vacuum dropout on recorder. Scans from 08:30 to 08:58 missed, scans from 10:00 to 10:30 missed, scans from 11:00 to 11:15 missed all due to tape problem. Tape MPIR0039 should be checked for wear! PI did not give adequate slew times for Azimuth unwraps. Tapes MPIR1239 and MPIR0039 shipped on 21/10/96 to Socorro via DHL.